OpenStudy (anonymous):

A man takes a step forward with a probability 0.4 and backwards with probability 0.6.Find the prob that at the end of 11 steps he is just 1 step away from the starting point?

OpenStudy (anonymous):

I know there will be 2 cases.. Case 1: He takes 6 steps forward and 5 steps backward Case 2: He takes 5 steps backward and 6 steps forward. I also know it is based on Bernoulli Trials,but i don't know how to put it mathematically.

OpenStudy (kropot72):

\[P(6\ of\ 11\ steps\ are\ forward)=\left(\begin{matrix}11 \\ 6\end{matrix}\right)\times0.4^{6}\times0.6^{5}\ .........(1)\] \[P(5\ of\ 11\ steps\ are\ forward)=\left(\begin{matrix}11 \\ 5\end{matrix}\right)\times0.4^{5}\times0.6^{6}\ .........(2)\] The values of probability calculated from equations (1) and (2) are added to find the required probability.

OpenStudy (anonymous):

you are taking forward as sucess? I mean how do u decide what is sucess and fail here

OpenStudy (kropot72):

For the man to finish one step away from the starting point after 11 steps means that there are two events that are a success: 1. The man takes a total of 6 steps forward and 5 steps backwards, resulting in him finishing one step in front of his starting point. 2. The man takes a total of 5 steps forward and 6 steps backwards, resulting in him finishing one step behind his starting point. The calculations that I have posted use the binomial distribution to find the values of probability for event 1 and event 2.